M7-19. Computing Bond Issue Price Abbington Inc. issues $700,000 of 9% bonds that pay interest semiannually and mature in 10 years. Com- pute the bond issu a. 8% per year compounded semiannually. b. 10% per year compounded semiannually.
Answer #1

a

Particulars Cash flow Discount factor Discounted cash flow
Interest payments-Annuity (4%,20 periods) 31,500.0 13.5903 428,095.28
Principle payments -Present value (4%,20 periods) 700,000 0.4564 319,470.86
A Bond price            747,566.14
Face value            700,000.00
Premium/(Discount)               47,566.14
Interest amount:
Face value 700,000
Coupon/stated Rate of interest 9.00%
Frequency of payment(once in) 6 months
B Interest amount 700000*0.09*6/12= 31500
Present value calculation:
yield to maturity/Effective rate 8.00%
Effective interest per period(i) 0.08*6/12= 4.000%
Number of periods:
Ref Particulars Amount
a Number of interest payments in a year                                     2
b Years to maturiy                                10.0
c=a*b Number of periods                                   20

b

Particulars Cash flow Discount factor Discounted cash flow
Interest payments-Annuity (5%,20 periods) 31,500.0 12.4622 392,559.63
Principle payments -Present value (5%,20 periods) 700,000 0.3769 263,822.64
A Bond price            656,382.26
Face value            700,000.00
Premium/(Discount)             -43,617.74
Interest amount:
Face value 700,000
Coupon/stated Rate of interest 9.00%
Frequency of payment(once in) 6 months
B Interest amount 700000*0.09*6/12= 31500
Present value calculation:
yield to maturity/Effective rate 10.00%
Effective interest per period(i) 0.1*6/12= 5.000%
Number of periods:
Ref Particulars Amount
a Number of interest payments in a year                                     2
b Years to maturiy                                10.0
c=a*b Number of periods                                   20